Parth (parthkohli):

I will state the question exactly as it is:

Parth (parthkohli):

If \(x,y \in [0, \pi/2)\):\[\sin^4 x + \cos^4 y + 2 = 4\sin x \cos y\]Then find the value of:\[\sin x + \sin y \]... The answer is given as 2.

OpenStudy (anonymous):

@Mehek14 @pinkbubbles

Parth (parthkohli):

I have two issues with this: one, \(\pi/2\) is not included. If I acknowledge this as a small mistake, even then, if \(\sin x + \sin y = 2\), then \(x = y = \pi/2\). But of course this does not satisfy the condition. Here is what *does* satisfy the condition: \(x = 0;~ y = \pi/2\)

Parth (parthkohli):

@ganeshie8

Parth (parthkohli):

@ikram002p @dan815

OpenStudy (anonymous):

@SolomonZelman

Parth (parthkohli):

@oldrin.bataku of course.

Parth (parthkohli):

Oops, I meant that \(y = 0; x = \pi/2\) satisfies the condition. Moving on...

imqwerty (imqwerty):

Pi/2 MUST be included.

imqwerty (imqwerty):

I've solved the question thoroughly nd m sure that pi/2 mst be included.

Parth (parthkohli):

Even then, the answer couldn't be 2, right?

Parth (parthkohli):

Also, how did you solve this question thoroughly if you don't mind explaining?

imqwerty (imqwerty):

K wait lemme switch on my laptop :)

Parth (parthkohli):

Did you solve this my maxima-minima or ...?

imqwerty (imqwerty):

No i jst simplified the expression

Parth (parthkohli):

Oh, great.

Parth (parthkohli):

Yeah, that's what I got too... finally.

Parth (parthkohli):

Thank you!

Parth (parthkohli):

So the answer isn't 2, it's 1, right?

Parth (parthkohli):

The question asked for sin x + sin y...

imqwerty (imqwerty):

yes its one :)

Parth (parthkohli):

Cool, I was wondering where the mistake was in my work. There wasn't any. I'll report this, thanks!

imqwerty (imqwerty):

ur welcome :)